OpenStudy (anonymous):

solve for X: log base 2 (2/X) = 3 + log base 2 X

OpenStudy (anonymous):

I get x=1/2. Your second line in the previous post doesn't follow. They are multiplied not added.

OpenStudy (anonymous):

how did you get 1/2 makaze?

OpenStudy (anonymous):

(why i can't give out medal)

OpenStudy (anonymous):

the answer is suppose to be x = 1/2

OpenStudy (anonymous):

I'll try to explain...

OpenStudy (anonymous):

I got 1/2

OpenStudy (anonymous):

sorry for the delay, but im a little bit tired

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

First you get log2(2)-log2(x)=3+log2(x), take the log2(2)=1 so you get -2=2log2(x) so after dividing by 2 you simply have 2^(-1)=x=1/2

OpenStudy (anonymous):

ok I have to go. hope you understand what I posted. I have to sleep a little bit.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

after i divide by 2 i get -1 = log2(x)

OpenStudy (anonymous):

and then 2^(-1) = x

OpenStudy (anonymous):

got it thank you